OpenStudy (anonymous):

Harriet rented an intermediate car for two days and drove it 500 miles. The daily rate was $36.99 and $0.32 per mile with the first 50 miles being free. If Harriet also had to pay 6 percent tax and $15 insurance per day, what was her total owed?

OpenStudy (anonymous):

Total number of miles chargable = 500 - 50 = 450 Total cost = (Daily cost x Number of days) + (Total number of miles chargable x Rate per mile) + (Insurance cost x Number of days) + (Tax Rate)

OpenStudy (anonymous):

Actually Total cost = ((Daily cost x Number of days) + (Total number of miles chargable x Rate per mile) + (Insurance cost x Number of days))(1+tax rate) So total cost = ((36.99*2)+(450*0.32)+(15*2))(1.06) = 262.8588 = $262.86